OpenStudy (carolina4567):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (carolina4567):

@phi

OpenStudy (carolina4567):

@iGreen

OpenStudy (carolina4567):

Please help me at least with first question

OpenStudy (igreen):

Quadratic form is: \(f(x) = ax^2 + bx + c\)

OpenStudy (igreen):

Wait..for here I think they're using: \(y = a(x - h)^2 + k\)

OpenStudy (igreen):

Let's just pick 3 for a, since it wants it to be between 4 and -4.

OpenStudy (igreen):

\(y = 3(x - 4)^2 + 12\)

OpenStudy (igreen):

I believe that's it.. Is your birthday in April or December..? @carolina4567

OpenStudy (igreen):

Okay, so that's correct. Your equation is: \(y = 3(x - 4)^2 + 12\).

OpenStudy (carolina4567):

So now we have to convert to standard form how would we do this ?

OpenStudy (igreen):

Yes.

OpenStudy (carolina4567):

standard form is y = a x^2 + b x + c where a, b and c are numbers

OpenStudy (igreen):

\(y = 3(x - 4)^2 + 12\) Standard form = \(y = ax^2 + bx + c\) \((a + b)^2 = a^2 + 2ab + b^2\) \((x - 4)^2 = x^2 + 2x(-4) - 4^2\) \((x - 4)^2 = x^2 - 8x- 16\) So: \(y = 3(x^2 - 8x - 16) + 12\) Distribute the 3: \(y = 3x^2 - 24x - 48 + 12\) Simplify: \(y = 3x^2 - 24x - 36\)

OpenStudy (carolina4567):

Oh so thats how u do it now i get it

OpenStudy (igreen):

Ugh, I don't know how but I got the equation wrong. It's supposed to be \(y = 3x^2 - 24x + 60\)

OpenStudy (igreen):

Oh, I see it now.

OpenStudy (igreen):

\(y=3(x−4)^2+12\) Standard form = \(y=ax^2+bx+c\) \((a+b)^2=a^2+2ab+b^2\) \((x−4)^2=x^2+2x(-4)+\color{red}4^2\) \((x−4)^2=x^2−8x\color{red}+16\) So: \(y=3(x^2−8x\color{red} +16)+12\) Distribute the 3: \(y=3x^2−24x\color{red} +48+12\) Simplify: \(y=3x^2−24x\color{red}{+60}\)

OpenStudy (igreen):

Okay, that's it. If you see these: '�', then click on a different question and click back. It'll be normal.

OpenStudy (carolina4567):

Oh cool yea it looks fine now

OpenStudy (igreen):

Okay, that's it..

OpenStudy (igreen):

You can graph your equation here: https://www.desmos.com/calculator Just copy and paste your equation in one of the boxes on the left.

OpenStudy (igreen):

Be sure to put in the exponents.

OpenStudy (carolina4567):

Yes i did

OpenStudy (igreen):

Do you see the graph?

OpenStudy (carolina4567):

Yes

OpenStudy (igreen):

Okay, the formula for the average rate of change is \(m = \dfrac{y_2-y_1}{x_2-x_1}\). Your graph will tell you that when x is 4, y is 12.(4, 12) And when x is 7, y is approximately 39.

OpenStudy (igreen):

Okay so our points are: \(y_2 \approx 39\) \(y_1 = 12\) \(x_2 = 7\) \(x_1 = 4\) Plug them in: \(m \approx \dfrac{39-12}{7-4}\) Subtract: \(m \approx \dfrac{27}{3}\) Divide: \(m \approx 9\) Remember, m is approximately 9, because 39 was approximate, and not exact.

OpenStudy (carolina4567):

Thank u so much for ur help @iGreen ur a life saver

OpenStudy (igreen):

No problem!